evaluate the following integral
∫ sqrt{(1 + cos 2x) / (1 – cos 2x)} dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \sqrt{\frac{1+cos2x}{1-cos2x}}dx
I = \int \sqrt{\frac{2cos^{2}x}{2sin^{2}x}}dx
I = \int \frac{cosx}{sinx}dx
sinx = t
cosxdx = dt
I = \int \frac{dt}{t}
I = log(t)+c
I = log(sinx)+c